# In-Context Imitation Learning via Next-Token Prediction

### v1

@@ 7 Limitations and Conclusion
This method has a few limitations. While results suggest that ICRT can generalize the primitive to unseen objects and certain primitives that resemble the ones in training (see Section 8.2.3), it is still unclear how to generalize to unseen primitives. Future works should investigate how scaling model capacity and scaling dataset can help with primitive-level generalization. In addition, ICRT assumes a fixed robot morphology with a fixed impedance controller. Future works can also investigate how to facilitate transfer between different robot morphologies by learning a unified policy on different robots. ICRT-Llama2 has a low inference frequency which may contribute to its low performance. We hope to speed up ICRT-Llama2 at inference time in the future.
@@
In summary, we present ICRT, where we study in-context imitation learning on a real robot. We do so by training a causal transformer model on sequences of robot trajectories, where trajectories of the same task are combined to serve as the context for performing the task. We also present a corresponding multi-task dataset to help facilitate this in-context learning. We find that by using robot sensorimotor trajectories as the context, the model can generalize the learned primitives to unseen objects and different environment configurations, especially in environments where more than one task is present.
Acknowledgments
@@
This research was performed at the AUTOLAB at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS ”People and Robots” (CPAR) Initiative. In their academic roles at UC Berkeley, Letian Fu, Huang Huang, Gaurav Datta, Lawrence Yunliang Chen, William Chung-Ho Panitch, Fangchen Liu, and Ken Goldberg are supported in part by donations from Autodesk, Meta, Google, Siemens, Toyota Research Institute, Bosch, and by equipment grants from PhotoNeo, Nvidia, NSF AI4OPT Centre, and Intuitive Surgical. L.Y. Chen is also supported by the National Science Foundation (NSF) Graduate Research Fellowship Program under Grant No. 2146752. We thank Xinyang Geng, Dantong Niu, and Chung Min Kim for their helpful discussions and feedback.

### v2

@@ VII Limitations and Conclusion
While results suggest that ICRT learns from the prompt trajectories and generalizes to unseen objects, tasks, and certain primitives that resemble the ones in training [^1] Please refer to the appendix and videos on the website for more detail., it is still unclear how to generalize to completely unseen action primitives. Future works should investigate how scaling model capacity and scaling dataset can help with primitive-level generalization. In addition, ICRT assumes a fixed robot morphology with a fixed impedance controller. Future works can also investigate how to facilitate transfer between different robot morphologies by learning a unified policy on different robots. ICRT-Llama2 has a low inference frequency which may contribute to its low performance. We hope to speed up ICRT-Llama2 at inference time in the future.
@@
In summary, we present ICRT, where we study in-context, multi-task imitation learning on a real robot. We achieve this by training a causal transformer model on sequences of robot trajectories, where trajectories from the same task are combined to provide context for task execution. Additionally, we introduce a multi-task dataset to facilitate this in-context learning approach. Our experiments show that by using robot sensorimotor trajectories as context, the model can generalize learned motion primitives to unseen objects and novel environment configurations, particularly in scenarios where multiple tasks are present.
VIII Acknowledgement
@@
This research was performed at the AUTOLAB at UC Berkeley in affiliation with the Berkeley AI Research (BAIR) Lab, and the CITRIS ”People and Robots” (CPAR) Initiative. In their academic roles at UC Berkeley, Letian Fu, Huang Huang, Gaurav Datta, Lawrence Yunliang Chen, William Chung-Ho Panitch, Fangchen Liu, and Ken Goldberg are supported in part by donations from Autodesk, Meta, Google, Siemens, Toyota Research Institute, Bosch, and by equipment grants from PhotoNeo, Nvidia, NSF AI4OPT Centre, and Intuitive Surgical. L.Y. Chen is also supported by the National Science Foundation (NSF) Graduate Research Fellowship Program under Grant No. 2146752. We thank Jiayi Pan, Xinyang Geng, Dantong Niu, and Chung Min Kim for their helpful discussions and feedback.
[…]
IX Supplementary Material
IX-A Scene Illustrations
We provide illustrations on the prompt trajectories and test scenes for the pick up the black dog and place in the pink bowl task in Figure 5. As mentioned in Section V, we collected 3 types of prompt trajectories and test ICRT on 5 tiers of scenes that are different from the scenes in the prompt trajectories.
Figure 5: Illustrations of the prompt trajectories (top) and test scenes (bottom) for the pick up the black dog and place in the pink bowl task. Three prompt trajectories of different types are collected. The test scenes are different from all prompt trajectories and 5 tiers of scenes with different number of distractors are considered.
IX-B Ablation Studies
In this section, we provide additional ablation experiments on a few core design choices and different prompting strategies.
IX-B1 Repeatability Experiments
We conduct experiments to evaluate the repeatability of the performance of ICRT. We conduct a pick up the black dog and place in the pink bowl task and a poke blue sponge task for 5 rollouts, where each rollout contains 5 trials as in Section V, resulting a total of 25 trials. We calculate the average and the standard deviation of the success rate. Results are shown in Table III. The low std from Table III suggests that the ICRT can reliably achieve the task.
Task Pick and Place Block Dog in Pink Bowl Poke Blue Sponge
Success Rate Ave. $\pm$ Std. 60% $\pm$ 0.5% 88% $\pm$ 3.2%
TABLE III: Repeatability experiments for a pick and place task and a poking task. Each task is conducted by 5 rollouts and each rollout contains 5 trials, resulting in a total of 25 trials.
Prompt Type No Distractor One Distractor Distractor Placement Two Prompts Three Prompts Success Rate 60% 80% 70% 80% 80%
TABLE IV: Experiments on different prompt types on a pick up black dog and place in the pink bowl task. The first three columns are results for a single prompt trajectory of different types, while the last two columns are that for using two and three prompts. Success rates are calculated over 5 trials for each experiment.
IX-B2 Prompt Trajectories
We conduct experiments on different prompt types to evaluate the effect of different prompt trajectories on task performance. We consider the task of picking up a black dog and placing in a pink bowl. We have three prompt trajectories of different types: one with no distractors, one with one distractor and one with one distractor placement, as shown in Appendix Figure 5 top. All three prompts trajectories are collected by human teleoprating the robot. The object locations and the placement locations at test time are different from that in all three prompts. As in Section V, for each prompt type, we conduct the task with 5 trials as shown in Appendix Figure 5 bottom. The average success rates are reported in Table IV. We conduct experiments with one prompt trajectory of different types (the first three columns in Table IV), two prompt trajectories and three prompt trajectories. All prompt types result in similar performance, indicating ICRT is not sensitive to the prompt trajectory types. We hypothesize this is because during the training, ICRT has seen different types and numbers of prompts.
IX-B3 Unseen Primitives
Task Grasp and Drop the Toy Tiger Grasp and Drop the Blue Sponge Put Blue Sponge to Right of Toy Tiger Success Rate 40% 80% 80%
TABLE V: Experiments on three tasks using two unseen primitives. Success rates are calculated over 5 trials for each experiment.
We evaluate the generalization capability of ICRT on primitives that are unseen during the training but resemble the training primitives. We consider two such unseen primitives: grasp and drop an object and put object A to the right of object B. We consider three tasks: grasp and drop a toy tiger, grasp and drop a blue sponge (unseen objects during training) and put the blue sponge to the right of the toy tiger. As in Section V, we conduct 5 tiers for each task. Experiment results are summarized in Table V, where ICRT shows decent success rate on all three tasks, suggesting that ICRT can generalize to some unseen primitives that resemble the training primitives.
IX-B4 Co-training
For training ICRT, we opt to separate the training into two stages: a pre-training phase where the model is pre-trained on the DROID dataset [khazatsky2024droid], and a fine-tuning phase where the model is trained on the ICIL-MT dataset. In this ablation, we experiment with whether these two can be combined into a single stage, where the policy is end-to-end trained with DROID and ICIL-MT. To balance the two datasets, we first calculate the median number of trajectories per task across the two datasets, then for each epoch, sample each task with the median number of trajectories. This allows each task to be equally represented in each epoch. We train the model for the same number of epochs as for ICRT fine-tuning and report the results in Table VI. The results indicate that the model does not converge as quickly in the combined stage and fails to respond to prompts and complete tasks effectively. We hypothesize two reasons for this: firstly, the dataset is heavily biased towards DROID, which contains 200 tasks compared to only 29 tasks in ICIL-MT, making it difficult for the model to learn the tasks as effectively as in the separate stage training. Future works can analyze the data mixture and how to train with large-scale datasets more effectively.
Pick and Place Poke Average
ICRT (Co-train) 13.3 ($\pm$5.8) 0.0 ($\pm$0.0) 6.7 ($\pm$3.0)
ICRT 65.0 ($\pm$7.3) 93.3 ($\pm$4.6) 79.2 ($\pm$4.6)
TABLE VI: Ablation on co-training with DROID [khazatsky2024droid]. Training both DROID and ICRT-MT datasets in a single stage leads to worse task performance in both pick-and-place and poking tasks. Same as Table I, we evaluated each task primitive using six tasks not seen during training, conducting five trials per task for a total of 30 trials per primitive and 60 trials overall to calculate average performance. For each model, we report the mean success rate for each task, the overall success rate, and the corresponding standard error in parentheses.
IX-C Detailed results
In this section, we present the per-task performance for the pick-and-place primitive (Table VII) and the poking primitive (Table VIII). For each action primitive, we design six unseen tasks (as defined in Section III), with three tasks evaluating in-domain object generalization (selected from yellow cube, red cube, black cube, pink bowl, and blue bear and three objects unseen during training (selected from radish, blue sponge, grey dog, and black dog).
Pick Object Place Location Yellow Cube Black Bowl Yellow Cube Grey Bowl Blue Bear Pink Bowl Radish Grey Bowl Black Dog Pink Bowl Blue Sponge Silver Pot Average Success ($\pm$ Std Err.) Goal Conditioned 40% 30% 20% 40% 40% 30% 33.3% ($\pm$6.5%) Octo 10% 0% 10% 10% 0% 0% 5.0% ($\pm$2.7%) OpenVLA 0% 0% 0% 50% 20% 0% 11.7% ($\pm$4.6%) ICRT-Llama2 40% 40% 40% 60% 40% 40% 43.3% ($\pm$7.9%) ICRT (DROID) 0% 0% 0% 0% 0% 0% 0.0% ($\pm$0.0%) ICRT (MT) 90% 50% 80% 90% 60% 90% 76.7% ($\pm$7.1%) ICRT +Prompt Loss 20% 10% 20% 40% 30% 10% 21.7% ($\pm$6.2%) ICRT (Co-train) 10% 0% 10% 0% 40% 20% 13.3% ($\pm$5.8%) ICRT 60% 50% 80% 50% 60% 90% 65.0% ($\pm$7.3%)
TABLE VII: Experimental results for the pick-and-place primitive. Here we list the 6 tasks that were evaluated and their corresponding success rate. For each model, we also report the mean success rate and the standard error.
Poke Object Radish Red Cube Grey Dog Black Cube Pink Bowl Blue Sponge Average Success ($\pm$ Std Err.) Goal Conditioned 0% 0% 0% 0% 40% 0% 6.7% ($\pm$4.6%) Octo 20% 0% 60% 0% 0% 0% 13.3% ($\pm$6.2%) OpenVLA 20% 0% 0% 0% 0% 0% 3.3% ($\pm$3.3%) ICRT-Llama2 60% 100% 80% 60% 60% 80% 73.3% ($\pm$8.2%) ICRT (DROID) 0% 0% 0% 0% 0% 0% 0.0% ($\pm$0.0%) ICRT (MT) 100% 100% 40% 60% 60% 60% 70.0% ($\pm$8.5%) ICRT +Prompt Loss 0% 20% 20% 80% 0% 20% 23.3% ($\pm$7.9%) ICRT (Co-train) 0% 0% 0% 0% 0% 0% 0.0% ($\pm$0.0%) ICRT 100% 100% 80% 80% 100% 100% 93.3% ($\pm$4.6%)
TABLE VIII: Experimental results for the poking primitive. Here we list the 6 tasks that were evaluated and their corresponding success rate. For each model, we also report the mean success rate and the standard error.
IX-D Hyperparameters
We provide the hyperparameters for both the pre-training and fine-tuning phase in Table IX and Table X.
Config Value
optimizer AdamW
base learning rate 1e-3
learning rate schedule cosine decay
batch size 64
weight decay 0.05
optimizer momentum $\beta_{1},\beta_{2}$ = 0.9, 0.999
warm up epoch 0.5
total epochs 4
proprioception noise 0.005
action noise 0
sequence length 512
brightness augmentation 0.1
contrast augmentation 0.2
num action prediction 16
TABLE IX: Pre-training Hyperparameters
Config Value
optimizer AdamW
base learning rate 5e-4
learning rate schedule cosine decay
batch size 64
weight decay 0.01
optimizer momentum $\beta_{1},\beta_{2}$ = 0.9, 0.999
warm up epoch 1.25
total epochs 125
proprioception noise 0.005
action noise 0
sequence length 512
brightness augmentation 0.1
contrast augmentation 0.2
num action prediction 16
TABLE X: Finetuning Hyperparameters
IX-E Parameterization
Proprioception The proprioception space is parameterized by the absolute end effector translation (x, y, z), a 6DoF rotation vector, and a continuous end-effector gripper state. This results in a 10-dimensional proprioception representation. The 6DoF rotation vector is flattened from the $\textit{SO}(3)$ rotation’s matrix’s first two rows.
Action We use delta end effector pose as the action parameterization. At each prediction step, the model predicts $t$ actions. Given absolute end effector action transforms in $T_{1},T_{2},\cdots,T_{t}$ in a trajectory and the current end-effector pose $T_{\text{ee}}$, we define the relative transforms that the model needs to predict as $T_{\text{ee}}^{-1}T_{1},T_{\text{ee}}^{-1}T_{2},\cdots T_{\text{ee}}^{-1}T_{t}$. We then append the continuous absolute gripper position to each delta action. Similar to proprioception, we present the delta action by the relative end effector translation and a 6DoF rotation. This results in a 10-dimensional action representation. When rolling out the predicted actions, in addition to temporal ensembling [zhao2023learning], we also use receding horizon control [diffusion], and select an action horizon of 10 steps.
IX-F System Information
All models are trained on 4 NVIDIA A100 80GB GPUs. ICRT pre-training on DROID takes 56 minutes and fine-tuneing on ICRT-MT takes 18 hours. ICRT-Llama7B takes roughly 28 hours to finetune. We report the inference speed of ICRT and ICRT-Llama2 in Table XI averaged over 100 steps. All tests are performed on a workstation with NVIDIA RTX 3090Ti and Intel i5-12400F with 64GB memory. We find that using the proposed formulation, which can leverage the KV cache, we can run ICRT-Llama2 at 10Hz naively.
Inference Frequency
ICRT 39.6 Hz
ICRT-Llama2 10.7 Hz
TABLE XI: Inference frequency of ICRT, averaged over 100 steps.